# Rational solutions of Painlevé V from Hankel determinants and the asymptotics of their pole locations

### v2

@@ 7 Conclusion
In this paper we have considered the particular scaling where we send $n\to\infty$ and rescale only the independent variable $t=ns$ (4.2). However, as evidenced also in Fig. 3, if ${\rho}$ or ${K}$ or both are large relative to $n$, the approximation needs to be modified.
If we let ${\rho}=n\varrho$, ${K}=n\varkappa$ then we would need to construct a different $g$–function where $\theta$ in (4.2) is replaced by
$\theta(z;s)=n\left(\frac{s}{z}+\varrho\ln\left(1-\frac{1}{z}\right)+\varkappa\ln z\right).$ (7.1)
Consequently the construction of the $g$–function, even under the one-cut assumption would significantly change. In practice this means that the set $EoT$ would have a different shape that depends on $\varrho,\varkappa$. While conceptually there is no major difference, we found that there are practical and significant obstacles in obtaining an effective description of the $g$–function under these assumptions. A separate analysis is needed but it is deferred to a future publication.
@@ Appendix A Airy parametrices
The complete construction of the approximation to the RHP of the main body of the paper requires the definition of a local solution to the final RHP which is known in the literature as an “Airy parametrix”. While this is quite standard, it may be useful for the reader to find here its complete and self contained definition. The origin of these definitions can be traced back to [26] but here we refer to App. A in [6]. We define
$A_{0}(\zeta)=\begin{bmatrix}\frac{{\mathrm{d}}}{{\mathrm{d}}\zeta}\textnormal{Ai}(\zeta)&{\rm e}^{i\frac{\pi}{3}}\frac{{\mathrm{d}}}{{\mathrm{d}}\zeta}\textnormal{Ai}\Big{(}{\rm e}^{-i\frac{2\pi}{3}}\zeta\Big{)}\\ \textnormal{Ai}(\zeta)&{\rm e}^{i\frac{\pi}{3}}\textnormal{Ai}\Big{(}{\rm e}^{-i\frac{2\pi}{3}}\zeta\Big{)}\\ \end{bmatrix},\hskip 14.22636pt\zeta\in\mathbb{C}$ (A.1)
where $\textnormal{Ai}(\zeta)$ is the Airy function, namely, a particular solution to Airy’s equation
$f(\zeta)^{\prime\prime}=\zeta f(\zeta)$
satisfying the following asymptotic behaviour as $\zeta\rightarrow\infty$ in the sector $-\pi<\textnormal{arg}\ \zeta<\pi$
$\textnormal{Ai}(\zeta)=\frac{\zeta^{-1/4}}{2\sqrt{\pi}}{\rm e}^{-\frac{2}{3}\zeta^{3/2}}\left(1-\frac{5}{48}\zeta^{-3/2}+\frac{385}{4608}\zeta^{-6/2}+\mathcal{O}\left(\zeta^{-9/2}\right)\right).$
We finally define the following piecewise-analytic matrix–valued function;
[…]
A0(ζ)[1101],arg ζ(2π3,0),
A0(ζ)[0111],arg ζ(π,2π3),
[…]
which solves the RHP with jumps for $\textnormal{arg}\,\zeta=-\pi,-\frac{2\pi}{3},0,\frac{2\pi}{3}$ as depicted in Fig. 11.
@@
The definition (A.2) is crafted in such a way that the following asymptotic expansion holds in any direction
${\bf A}(\zeta)=\frac{\zeta^{\sigma_{3}/4}}{2\sqrt{\pi}}\begin{bmatrix}-1&i\\ 1&i\\ \end{bmatrix}\Bigg{\{}\mathbf{1}+\frac{1}{48\zeta^{3/2}}\begin{bmatrix}1&6i\\ 6i&-1\\ \end{bmatrix}+\mathcal{O}\left(\zeta^{-6/2}\right)\Bigg{\}}{\rm e}^{-\frac{2}{3}\zeta^{3/2}\sigma_{3}}.$ (A.3)
@@ A.1 The Airy parametrix for a three main-arc intersection
The basic building block discussed in the previous section can be used to construct a more complicated local parametrix that needs to be used in a neighbourhood of a $3$–arc intersection, like the case of the point $z=b$ for the asymptotic in $EoT$ for $\mathrm{Re}\,(s)>0$, Sec 5.2.1.
@@
Consider the jump matrices of the RHP 5.7 on the arcs that originate or terminate at $z=b$ (see (5.56)); they are summarized in the left pane of Fig. 12. There, we have denoted by $\varphi_{0},\varphi_{1},\varphi_{2}$ the restrictions of $\varphi$ to the sectors in $\mathbb{D}_{b}$. These sectors are defined as follows: $\mathscr{S}_{0}$ bounded by the arcs $\Gamma_{b}^{1},\Gamma_{b}^{a_{+}}$, $\mathscr{S}_{1}$ bounded by the arcs $\Gamma_{b}^{a_{+}},\Gamma_{a_{-}}^{b}$ and $\mathscr{S}_{2}$ bounded by the arcs $\Gamma_{a_{-}}^{b},\Gamma_{b}^{1}$, respectively (see Fig. 8, left pane). They are related to each other as follows
$\varphi_{0}(z)+\varphi_{1}(z)=2\Omega_{2},\ \ z\in\Gamma_{b}^{a_{+}},$ (A.4)
$\varphi_{0}(z)+\varphi_{2}(z)=2\Omega_{1},\ \ z\in\Gamma_{b}^{1},$ (A.5)
$\varphi_{1}(z)+\varphi_{2}(z)=0,\ \ z\in\Gamma_{a_{-}}^{b}.$ (A.6)
Observing that $\varphi_{0}(b)=-\Omega_{1}-\Omega_{2}$ we can define the local conformal map $\zeta(z)$ by the expression.
$\frac{4}{3}\zeta^{\frac{3}{2}}=n\varphi_{0}(z)+n(\Omega_{1}+\Omega_{2}),\ \ z\in\mathscr{S}_{0},$ (A.7)
where the radical of $\zeta$ is intended in the sense of principal determination (with the branch cut extending along $\zeta\in{\mathbb{R}}_{-}$). If we perform the analytic extension of this definition to the sectors $\mathscr{S}_{1},\mathscr{S}_{2}$ we obtain the following relations;
$n\varphi_{1}(z)$ $=-n\varphi_{0}(z)+2\Omega_{2}=-\frac{4}{3}\zeta^{\frac{3}{2}}-n\Omega_{1}+n\Omega_{2},\ \ z\in\mathscr{S}_{1}$ (A.8)
$n\varphi_{2}(z)$ $=-n\varphi_{0}(z)+2\Omega_{1}=-\frac{4}{3}\zeta^{\frac{3}{2}}-n\Omega_{2}+n\Omega_{1},\ \ z\in\mathscr{S}_{2}.$ (A.9)
@@
Moreover we have denoted by $\widetilde{Q}(z)$ the analytic extension of $Q(z)$ in the full disk $\mathbb{D}_{b}$ from the upper part, so that
$\widetilde{Q}(z)=\left\{\begin{array}[]{cc}Q(z)&\Big{\{}z:\ \mathrm{Im}\,\zeta(z)>0\Big{\}}\cap\mathbb{D}_{b}\\[5.0pt] {\rm e}^{2i\pi{{\rho}}}Q(z)&\Big{\{}z:\ \mathrm{Im}\,\zeta(z)<0\Big{\}}\cap\mathbb{D}_{b}.\end{array}\right.$ (A.10)
With these notation in place and recalling the relationship between $\varkappa$ and ${\rm e}^{2i\pi{{\rho}}}$ (5.39), the jump matrices are as indicated in Fig. 12, left pane; note that we have also re-oriented some of the arcs; note that the jump matrix on the left ray is the result of the multiplication of the jumps on the two arcs $\mathscr{L}_{+}^{(a_{-},b)},\mathscr{L}_{+}^{(b,a_{+})}$.
Figure 12: The local RHP 5.7 in $\mathbb{D}_{b}$ (left pane) and its transformation to the $\zeta$ plane.
With these preparations, the solution of the RHP with the jump matrices indicated in the right pane of Fig. 12 is written as follows in terms of the matrix ${\bf A}(z)$ defined in (A.2):
$\widetilde{\bf P}(\zeta)=\left\{\begin{array}[]{ll}{\bf A}(\zeta){\rm e}^{\left(\frac{2}{3}\zeta^{\frac{3}{2}}+\frac{\Omega_{1}-\Omega_{2}}{2}\right)\sigma_{3}}\left(\frac{\varkappa}{(\varkappa+1)\widetilde{Q}(z)}\right)^{\frac{\sigma_{3}}{2}}{\rm e}^{\frac{i\pi\sigma_{3}}{2}}&\arg(\zeta)\in\left(\frac{\pi}{3},\pi\right)\\ {\bf A}(\zeta){\rm e}^{\left(\frac{2}{3}\zeta^{\frac{3}{2}}+\frac{\Omega_{2}-\Omega_{1}}{2}\right)\sigma_{3}}\left(\frac{1}{\varkappa(\varkappa+1)\widetilde{Q}(z)}\right)^{\frac{\sigma_{3}}{2}}{\rm e}^{-\frac{i\pi\sigma_{3}}{2}}&\arg(\zeta)\in\left(-\pi,-\frac{\pi}{3}\right)\\ {\bf A}(\zeta){\rm e}^{\left(\frac{2}{3}\zeta^{\frac{3}{2}}+\frac{\Omega_{2}+\Omega_{1}}{2}\right)\sigma_{3}}\begin{bmatrix}0&-\widetilde{Q}(z)\\ \frac{1}{\widetilde{Q}(z)}&0\end{bmatrix}\left(\frac{\varkappa+1}{\varkappa\widetilde{Q}(z)}\right)^{\frac{\sigma_{3}}{2}}{\rm e}^{\frac{-i\pi\sigma_{3}}{2}}&\arg(\zeta)\in\left(-\frac{\pi}{3},\frac{\pi}{3}\right).\end{array}\right.$ (A.14)

### v3

@@ 7 Conclusion
In this paper, we have considered the particular scaling where we send $n\to\infty$ and rescale only the independent variable $t=ns$ (4.1). However, as evidenced also in Figure 3, if $\rho$} or K{K} or both are large relative to $n$, the approximation needs to be modified. If we let ${\rho}=n\varrho$, ${K}=n\varkappa$, then we would need to construct a different $g$-function where $\theta$ in (4.1) is replaced by
$\theta(z;s)=n\left(\frac{s}{z}+\varrho\ln\left(1-\frac{1}{z}\right)+\varkappa\ln z\right).$
Consequently, the construction of the $g$-function, even under the one-cut assumption would significantly change. In practice, this means that the set $\rm EoT$} would have a different shape that depends on $\varrho$, $\varkappa$. While conceptually there is no major difference, we found that there are practical and significant obstacles in obtaining an effective description of the $g$-function under these assumptions. A separate analysis is needed but it is deferred to a future publication.
@@ Appendix A Airy parametrices
The complete construction of the approximation to the RHP of the main body of the paper requires the definition of a local solution to the final RHP which is known in the literature as an “Airy parametrix”. While this is quite standard, it may be useful for the reader to find here its complete and self contained definition. The origin of these definitions can be traced back to [26] but here we refer to [6, Appendix A]. We define
$A_{0}(\zeta)=\begin{bmatrix}\displaystyle\frac{{\mathrm{d}}}{{\mathrm{d}}\zeta}\textnormal{Ai}(\zeta)&\displaystyle{\rm e}^{{\rm i}\frac{\pi}{3}}\frac{{\mathrm{d}}}{{\mathrm{d}}\zeta}\textnormal{Ai}\bigl({\rm e}^{-{\rm i}\frac{2\pi}{3}}\zeta\bigr)\\ \textnormal{Ai}(\zeta)&\displaystyle{\rm e}^{{\rm i}\frac{\pi}{3}}\textnormal{Ai}\bigl({\rm e}^{-{\rm i}\frac{2\pi}{3}}\zeta\bigr)\\ \end{bmatrix},\qquad\zeta\in\mathbb{C},\vskip-2.84526pt$
where $\textnormal{Ai}(\zeta)$ is the Airy function, namely, a particular solution to Airy’s equation $f(\zeta)^{\prime\prime}=\zeta f(\zeta)$ satisfying the following asymptotic behaviour as $\zeta\rightarrow\infty$ in the sector $\pi<\operatorname{arg}\zeta<\pi$:
$\textnormal{Ai}(\zeta)=\frac{\zeta^{-1/4}}{2\sqrt{\pi}}{\rm e}^{-\frac{2}{3}\zeta^{3/2}}\left(1-\frac{5}{48}\zeta^{-3/2}+\frac{385}{4608}\zeta^{-6/2}+\mathcal{O}\bigl(\zeta^{-9/2}\bigr)\right).\vskip-2.84526pt$
We finally define the following piecewise-analytic matrix-valued function:
${\bf A}(\zeta)=\begin{cases}A_{0}(\zeta),&\operatorname{arg}\zeta\in\bigl(0,\frac{2\pi}{3}\bigr),\\ A_{0}(\zeta)\begin{bmatrix}1&0\\ -1&1\\ \end{bmatrix},&\operatorname{arg}\zeta\in\bigl(\frac{2\pi}{3},\pi\bigr),\\[10.0pt] A_{0}(\zeta)\begin{bmatrix}1&-1\\ 0&1\\ \end{bmatrix}&\operatorname{arg}\zeta\in\bigl(-\frac{2\pi}{3},0\bigr),\\[10.0pt] A_{0}(\zeta)\begin{bmatrix}0&-1\\ 1&1\\ \end{bmatrix},&\operatorname{arg}\zeta\in\bigl(-\pi,-\frac{2\pi}{3}\bigr),\end{cases}$ (A.1)
which solves the RHP with jumps for $\operatorname{arg}\zeta=-\pi,-\frac{2\pi}{3},0,\frac{2\pi}{3}$ as depicted in Figure 11.
@@
The definition (A.1) is crafted in such a way that the following asymptotic expansion holds in any direction:
${\bf A}(\zeta)=\frac{\zeta^{\sigma_{3}/4}}{2\sqrt{\pi}}\begin{bmatrix}-1&{\rm i}\\ 1&{\rm i}\\ \end{bmatrix}\left\{\mathbf{1}+\frac{1}{48\zeta^{3/2}}\begin{bmatrix}1&6{\rm i}\\ 6{\rm i}&-1\\ \end{bmatrix}+\mathcal{O}\bigl(\zeta^{-6/2}\bigr)\right\}{\rm e}^{-\frac{2}{3}\zeta^{3/2}\sigma_{3}}.$
@@ A.1 The Airy parametrix for a three main-arc intersection
The basic building block discussed in the previous section can be used to construct a more complicated local parametrix that needs to be used in a neighbourhood of a $3$-arc intersection, like the case of the point $z=b$ for the asymptotic in $\rm EoT$} for $\operatorname{Re}(s)>0$, Section 5.2.1.
@@
Consider the jump matrices of the Riemann–Hilbert Problem 5.12 on the arcs that originate or terminate at z=b{z=b} (see (5.17)). They are summarized in the left pane of Figure 12. There, we have denoted by $\varphi_{0}$$\varphi_{1}$$\varphi_{2}$ the restrictions of $\varphi$ to the sectors in $\mathbb{D}_{b}$. These sectors are defined as follows: $\mathscr{S}_{0}$ bounded by the arcs $\Gamma_{b}^{1}$, $\Gamma_{b}^{a_{+}}$, $\mathscr{S}_{1}$ bounded by the arcs $\Gamma_{b}^{a_{+}}$, $\Gamma_{a_{-}}^{b}$ and $\mathscr{S}_{2}$ bounded by the arcs $\Gamma_{a_{-}}^{b}$$\Gamma_{b}^{1}$, respectively (see Figure 8, left pane). They are related to each other as follows:
$\varphi_{0}(z)+\varphi_{1}(z)=2\Omega_{2},\qquad z\in\Gamma_{b}^{a_{+}},$
$\varphi_{0}(z)+\varphi_{2}(z)=2\Omega_{1},\qquad z\in\Gamma_{b}^{1},$
$\varphi_{1}(z)+\varphi_{2}(z)=0,\qquad z\in\Gamma_{a_{-}}^{b}.$
Observing that $\varphi_{0}(b)=-\Omega_{1}-\Omega_{2}$, we can define the local conformal map $\zeta(z)$ by the expression
$\frac{4}{3}\zeta^{\frac{3}{2}}=n\varphi_{0}(z)+n(\Omega_{1}+\Omega_{2}),\qquad z\in\mathscr{S}_{0},$
where the radical of $\zeta$ is intended in the sense of principal determination (with the branch cut extending along $\zeta\in{\mathbb{R}}_{-}$). If we perform the analytic extension of this definition to the sectors $\mathscr{S}_{1}$, $\mathscr{S}_{2}$, we obtain the following relations:
$n\varphi_{1}(z)=-n\varphi_{0}(z)+2\Omega_{2}=-\frac{4}{3}\zeta^{\frac{3}{2}}-n\Omega_{1}+n\Omega_{2},\qquad z\in\mathscr{S}_{1},$
$n\varphi_{2}(z)=-n\varphi_{0}(z)+2\Omega_{1}=-\frac{4}{3}\zeta^{\frac{3}{2}}-n\Omega_{2}+n\Omega_{1},\qquad z\in\mathscr{S}_{2}.$
@@
Moreover, we have denoted by $\widetilde{Q}(z)$ the analytic extension of $Q(z)$ in the full disk $\mathbb{D}_{b}$ from the upper part, so that
$\widetilde{Q}(z)=\begin{cases}Q(z),&\{z\mid\operatorname{Im}\zeta(z)>0\}\cap\mathbb{D}_{b},\\ {\rm e}^{2{\rm i}\pi{{\rho}}}Q(z),&\{z\mid\operatorname{Im}\zeta(z)<0\}\cap\mathbb{D}_{b}.\end{cases}$
With these notation in place and recalling the relationship between $\varkappa$ and ${\rm e}^{2{\rm i}\pi{{\rho}}}$ (5.10), the jump matrices are as indicated in Figure 12, left pane. Note that we have also re-oriented some of the arcs. Note that the jump matrix on the left ray is the result of the multiplication of the jumps on the two arcs $\mathscr{L}_{+}^{(a_{-},b)}$, $\mathscr{L}_{+}^{(b,a_{+})}$.
Figure 12: The local Riemann–Hilbert Problem 5.12 in $\mathbb{D}_{b}$ (left pane) and its transformation to the $\zeta$ plane.
With these preparations, the solution of the RHP with the jump matrices indicated in the right pane of Figure 12 is written as follows in terms of the matrix ${\bf A}(z)$ defined in (A.1):
$\widetilde{\bf P}(\zeta)=\begin{cases}\displaystyle{\bf A}(\zeta){\rm e}^{(\frac{2}{3}\zeta^{\frac{3}{2}}+\frac{\Omega_{1}-\Omega_{2}}{2})\sigma_{3}}\left(\frac{\varkappa}{(\varkappa+1)\widetilde{Q}(z)}\right)^{\frac{\sigma_{3}}{2}}{\rm e}^{\frac{{\rm i}\pi\sigma_{3}}{2}},&\arg(\zeta)\in\bigl(\frac{\pi}{3},\pi\bigr),\\ \displaystyle{\bf A}(\zeta){\rm e}^{(\frac{2}{3}\zeta^{\frac{3}{2}}+\frac{\Omega_{2}-\Omega_{1}}{2})\sigma_{3}}\left(\frac{1}{\varkappa(\varkappa+1)\widetilde{Q}(z)}\right)^{\frac{\sigma_{3}}{2}}{\rm e}^{-\frac{{\rm i}\pi\sigma_{3}}{2}},&\arg(\zeta)\in\bigl(-\pi,-\frac{\pi}{3}\bigr),\\ \displaystyle{\bf A}(\zeta){\rm e}^{(\frac{2}{3}\zeta^{\frac{3}{2}}+\frac{\Omega_{2}+\Omega_{1}}{2})\sigma_{3}}\begin{bmatrix}0&-\widetilde{Q}(z)\\ \displaystyle\frac{1}{\widetilde{Q}(z)}&0\end{bmatrix}\left(\frac{\varkappa+1}{\varkappa\widetilde{Q}(z)}\right)^{\frac{\sigma_{3}}{2}}{\rm e}^{\frac{-{\rm i}\pi\sigma_{3}}{2}},&\arg(\zeta)\in\bigl(-\frac{\pi}{3},\frac{\pi}{3}\bigr).\end{cases}$
Acknowledgements
The second author completed the work during his tenure as Royal Society Wolfson Visiting Fellow (RSWVF/R2/242024) at the School of Mathematics in Bristol University. The work was supported in part by the Natural Sciences and Engineering Research Council of Canada (NSERC) grant RGPIN-2023-04747. Both authors thank the anonymous referees for the detailed reports and bibliographic improvements.